OpenStudy (anonymous):

a machine gun fires 25gm bullets at the rate of 600bullets per min with a speed of 200m/s. calculate the force required to keep the gun in position.

OpenStudy (anonymous):

plzz solve...

OpenStudy (anonymous):

600=N, m=25*10^-3, M=N*m=15kg. P=M*V; P=15*200=3000 kg*m/s; F=P/t; F=3000/60=50 n.